OpenStudy (anonymous):

in the case of the washing macine why is the acceleration so high?..its hard to imagine such a big number for its acceleration when the velocity is so much lower...

OpenStudy (anonymous):

it's angular acceleration on the other hand acceleration is high as to keep the whirl motion so that clothes may be washed thoroughly. its again like after a certain number of rotations the direction of current is reversed thus the rotation of machine reverses opposing the momentum of water. you need a heavy motor to do such a task. a small motor may not take such a load.

OpenStudy (ganpat):

Simple language, just to create more turbulence.. and the motor is design for low speed high torque, to take the load inside the machine..

OpenStudy (kainui):

Acceleration is a change in velocity. But remember velocity is a vector so it has a speed AND direction. So an acceleration doesn't always mean that you're changing speed. An acceleration in this case is a change in direction, and since it's going in a circle and not a straight line it is constantly accelerating by changing direction!

OpenStudy (anonymous):

im still not getting how changing direction adds metres per second...is it something to do with how far on the circumference the radius line covers in a second?

OpenStudy (anonymous):

the area it covers travelling in one second..

OpenStudy (anonymous):

think about this, follow a single point along the circumference of the washing machines drum... at any time (lets say at time t= 0) that point has a velocity vector in the x and y directions that are tangential to the circumference ( and a z component if your washing machine is rocking). A fraction of a second later (t+dt) that point has moved further along the circumference. its tangential velocity has different x and y coordinates. we know that acceleration is dv/dt (change in velocity divided by change in time, remember velocity is a vector) an example. V(at t=0sec) = <5i+2j> V(at=5sec)=<2i+5> the velocities magnitude are the same, |V| = Squareroot(29) a=dv/dt dv = V(5) - V(0) = <-3i+3j> [m/s] dt = 5-0 = 5 [seconds] dv/dt =<-3i+3j> / 5 seconds =squareroot(18)/5 [m/s^2] =0.848 m/s^2